OpenStudy (anonymous):

a water tank is 6 3/4 m long, 5 1/2 m wide and 6 2/3 m high. how much water will it contain when 5/8 full?

OpenStudy (anonymous):

first calculate its volume V=3/4 * 5 1/2 * 6 2/3

OpenStudy (anonymous):

\[V = \frac{ 3 }{ 4 } \times5\frac{ 1 }{ 2 }\times 6 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

sorry it is 6 3/4

OpenStudy (anonymous):

volume is 0.75*5.5*6.666 = 27.5 5/8 th part of 27.5 is 17.1875

OpenStudy (anonymous):

v = 27.84 m^3 volume of water = 5/8 v

OpenStudy (anonymous):

@Simonbane is thats answer

OpenStudy (anonymous):

did you change 3/4 into 6 3/4?

OpenStudy (anonymous):

6.75 * 5.5 * 6.67 = 247.62375

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

divided by 0.625?

OpenStudy (anonymous):

i am sorry.............6 2/3 to 6 3/4

OpenStudy (anonymous):

no. this is the given 6 3/4 long, 5 1/2 wide and 6 2/3 high

OpenStudy (anonymous):

ok then V = 247.5

OpenStudy (anonymous):

247.62375 or 247. 624

OpenStudy (anonymous):

then what should i do next? subtract it by 0.625 or divide it by 0.625? @uzumakhi

OpenStudy (anonymous):

you should divide it

OpenStudy (anonymous):

so my answer is 396. 198? @uzumakhi

OpenStudy (anonymous):

try to multiple it